Q. GE is the only producer of new jet engines for general aviation aircraft. Demand for a single engine is P = 2,000,000 – Q while the MCs of producing an engine are: MC = 1,999 Q.

a) What would be the monopoly price and quantity of these engines?

b) What economic profit would GE earn on the sale of these engines?

c) What would happen to price and quantity if the market were competitive (assuming the same costs)?

What if we have Q = 25 – 0.5P and TC = 50 + 2Q?
Answer #1

(Question 1)

(a)

Monopolist will equate MR and MC.

TR = P x Q = 2,000,000Q - Q2

MR = dTR/dQ = 2,000,000 - 2Q

2,000,000 - 2Q = 1,999Q

2,001Q = 2,000,000

Q = 999.5

P = 2,000,000 - 999.5 = 1,999,000.5

(b)

MC = 1,999 x 999.5 = 1,998,000.5

Profit = Q x (P - MC) = 999.5 x (1,999,000.5 - 1,998,000.5) = 999.5 x 1,000 = 999,500

(c)

In perfect competition, P = MC.

2,000,000 - Q = 1,999Q

2,000Q = 2,000,000

Q = 1,000

P = MC = 1,999 x 1,000 = 1,999,000

NOTE: As HOMEWORKLIB Answering Policy, 1st question has been answered.